OpenStudy (haleyelizabeth2017):

What is the standard form of the equation with rho=18, phi=(5pi)/4? Please no direct answers I have gotten to: x cos phi + y sin phi - p=0 x cos (5 pi/4) + y sin (5 pi/4) - 18 = 0 But I have no idea on where to go from here.

OpenStudy (shaleiah):

@Nnesha

OpenStudy (shaleiah):

@Directrix

OpenStudy (xapproachesinfinity):

you rectangular form?

OpenStudy (xapproachesinfinity):

the very fact that you talking about rho=18 i think you were essentially looking for a rectangular equation

OpenStudy (xapproachesinfinity):

converting from spherical to rectangular x^2+y2+z^2=rph^2

OpenStudy (xapproachesinfinity):

rho^2

OpenStudy (xapproachesinfinity):

hey talk to me! haha so i know what exactly you want to find

OpenStudy (haleyelizabeth2017):

Sorry haha was afk

OpenStudy (xapproachesinfinity):

i see this is not conversion anyway

OpenStudy (haleyelizabeth2017):

I have no idea...the information ^^ up there was all I was given.

OpenStudy (xapproachesinfinity):

so the standard form is given by xcos(phi)+ysin(phi)-p=0

OpenStudy (xapproachesinfinity):

which you already did xcos(5pi/4)+ysin(5pi/4)-18=0

OpenStudy (haleyelizabeth2017):

yeah...

OpenStudy (xapproachesinfinity):

now you just need to evaluate cos(5pi/4) and sin(5pi/4)

OpenStudy (xapproachesinfinity):

you should be familiar with trigonometry, so what the problem here?

OpenStudy (haleyelizabeth2017):

One second..

OpenStudy (xapproachesinfinity):

cos(5pi/4)=?

OpenStudy (haleyelizabeth2017):

.998

OpenStudy (haleyelizabeth2017):

sin(5pi/4)=.068

OpenStudy (xapproachesinfinity):

don't use calculator! if you know trig you know that \[\cos (\pi/4) =\sqrt{2}/2\]

OpenStudy (haleyelizabeth2017):

-_-

OpenStudy (xapproachesinfinity):

but we are looking for 5pi/4 so you need to know what quadrant is the angle 5pi/4 in

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

that angle is in the 3th quadrant so cos there is negative then cos(5pi/4)=-sqrt(2)/2

OpenStudy (xapproachesinfinity):

with the same strategy sin(5pi/4)=-sqrt(2)/2 so the equation becomes \[-\frac{\sqrt{2}}{2 }x-\frac{\sqrt{2}}{2}y=18\]

OpenStudy (haleyelizabeth2017):

Sorry